# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 12.1 Variations
The ${\mathcal{X}}_{p}$ may depend on the choice of the total order $\prec$ on $W$. Therefore it is not clear that ${\mathcal{X}}_{p}$ is perpendicular to ${\mathcal{X}}_{q}$ when $\ell(p)=\ell(q)$. In fact computer assisted computations in ${\sf K}_{T}(G/B)$ indicate that this fails already for type ${\bf B}_{3}$ with $\ell(p)=\ell(q)=3$ and for type ${\bf C}_{3}$ with $\ell(p)=\ell(q)=2$.
@@ 13 Generalised flag varieties
Let $P$ be a parabolic subgroup containing $B$. (If ${\sf k}={\mathbb{Z}}$, then $P$ means $\mathbb{P}$.) We seek a full exceptional collection on $G/P$. Let $W_{P}$ be the Weyl group of $P$, generated inside $W$ by the $s_{\alpha}$ with $\alpha$ simple and $P_{\alpha}\subset P$. Let $W^{P}$ be the set of minimal coset representatives of $W/W_{P}$, cf.  Lemma 4.5. According to Steinberg [44] we may use as generators of the $R(G)$-module $R(P)$, over ${\mathbb{C}}$, the classes of the irreducible $P$-modules with highest weight $e_{v}$ where $v$ runs over $W^{P}$, and $e_{v}$ is still defined as in 3.2. Inspired by that, we restrict our total order $\prec$ from $W$ to $W^{P}$.
@@
The map $W\to W^{P}$ which sends $w\in W$ to the minimal representative of the coset $wW_{P}$ is a poset map. If one has chosen a total order on $W^{P}$ refining the (restriction to $W^{P}$) of the Bruhat order, then this chosen order can be extended to a total order on $W$ that refines the Bruhat order on $W$. We already had to choose $\prec$ on $W$ in section 4.2, so it makes sense to keep that order $\prec$ and restrict it to $W^{P}$.
For $v\in W^{P}$ we will find ${\hat{X}}_{v}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ such that
@@
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$. Let $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ and $v,w\in W^{P}$.
@@
    $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes{\hat{X}}_{v},N\otimes{\hat{X}}_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(M,N)$,
@@
For $v\in W^{P}$, denote ${\hat{\sf X}}_{v}$ to be the strictly full subcategory of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ generated by $\{M\otimes{\hat{X}}_{v}\mid M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\}$. Then the category $\mathcal{D}$ has a $G$-linear semiorthogonal decomposition
$\mathcal{D}=\langle{\hat{\sf X}}_{v}\rangle_{v\in W^{P}}$ (45)
with respect to the order $\prec$ on $W^{P}$. Each subcategory ${\hat{\sf X}}_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
Definition 23.
We put $\hat{\mathcal{X}}_{v}:=\mathcal{L}_{G/P}({\hat{X}}_{v})$ for $v\in W^{P}$.
@@
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$. Let $v,w\in W^{P}$.
@@
    The triangulated hull of $\{\hat{\mathcal{X}}_{v}\mid v\in W^{P}\}$ is $\mathcal{D}$.
In other words, the collection of objects $(\hat{\mathcal{X}}_{v})_{v\in W}$ is a full exceptional collection in $\mathcal{D}$.
To prove these Theorems we will need to find replacements–often with hatted notation–of the key ingredients used in the case $P=B$. This will take the rest of the section.
Recall that by [29, II Corollary 4.7] we may view ${\mathrm{rep}}(G)$ as a subcategory of ${\mathrm{rep}}(P)$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ as a subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$, so that we may suppress ${\mathrm{res}}^{G}_{P}$ in notations. Similarly we may suppress ${\mathrm{res}}^{P}_{B}$.
@@
The strategy for constructing $\hat{\mathcal{X}}_{v}$ is the same as for ${\mathcal{X}}_{v}$, but we do not know the precise relation between ${\mathcal{X}}_{v}$, $\hat{\mathcal{X}}_{v}$ and $G/B\to G/P$. More specifically, we do not know if ${\mathcal{X}}_{v}$ is always the pull back of $\hat{\mathcal{X}}_{v}$ for $v\in W^{P}$.
@@
Let $\lambda\in X(T)$, $\alpha\in\Pi$. Then $s_{\alpha}\lambda\leq_{e}\lambda$ if and only if $s_{\alpha}\lambda\geq_{d}\lambda$.
@@
Write $\lambda=w\nu$ with $\nu$ dominant and $w$ minimal. We have $s_{\alpha}w>w$ or $s_{\alpha}w<w$. If $s_{\alpha}w>w$, then $\lambda$ is a weight of $P(s_{\alpha}\lambda)$ and the socle of $P(s_{\alpha}\lambda)$, which has weight $s_{\alpha}\lambda$, must lie in the $B$-submodule generated by the weight space of weight $\lambda$. This implies $s_{\alpha}\lambda\leq_{d}\lambda$. The other case is similar. ∎
Notations 13.7.
Let $\Pi^{P}$ be the set of simple roots $\alpha$ for which $P_{\alpha}\subseteq P$. Let $\Pi^{\notin P}$ be the set of simple roots outside $\Pi^{P}$. We say that $\lambda$ is $P$-dominant if $(\alpha^{\vee},\lambda)\geq 0$ for all $\alpha\in\Pi^{P}$. The set of $P$-dominant weights is denoted $X^{P}_{+}$. Every $W_{P}$-orbit of weights intersects $X^{P}_{+}$ in a unique element. Let $G_{P}$ be the semisimple subgroup of $P$ with $\Pi^{P}$ as simple roots. Note that $G_{P}$ is simply connected.
@@
Let $v\in W^{P}$. Then $P(-e_{v})$ is a $P$-module and therefore $P(-e_{v})^{*}$ is a $P$-module. Conversely, if $e_{w}$ is $P$-dominant, then $w\in W^{P}$.
@@
By [47, Proposition 2.2.5], there is a sequence $\alpha_{1},\cdots,\alpha_{n}$ with $\alpha_{i}\in\Pi^{P}$ and ${\mathrm{ind}}_{B}^{P}={\mathrm{ind}}_{B}^{P_{\alpha_{1}}}\circ\cdots\circ{\mathrm{ind}}_{B}^{P_{\alpha_{n}}}$. It therefore suffices to see that ${\mathrm{ind}}_{B}^{P_{\alpha}}(P(-e_{v}))=P(-e_{v})$ for $\alpha\in\Pi^{P}$. Now $v$ is a minimal coset representative in $W/W_{P}$, so $vs_{\alpha}>v$ for $\alpha\in\Pi^{P}$. Then $s_{\alpha}v^{-1}>v^{-1}$, so $s_{\alpha}v^{-1}w_{0}<v^{-1}w_{0}$, and $s_{\alpha}\star v^{-1}w_{0}=v^{-1}w_{0}$. Thus ${\mathrm{ind}}_{B}^{P_{\alpha}}(P(-e_{v}))={\sf H}_{s_{\alpha}}{\sf H}_{v^{-1}w_{0}}(-w_{0}ve_{v})={\sf H}_{v^{-1}w_{0}}(-w_{0}ve_{v})=P(-e_{v})$.
Conversely, let $e_{w}$ be $P$-dominant and $\alpha\in\Pi^{P}$. Now $we_{w}$ is dominant and $w^{-1}$ is minimal amongst the $z$ with $zwe_{w}=e_{w}$. Thus if $s_{\alpha}e_{w}=e_{w}$ then $s_{\alpha}w^{-1}>w^{-1}$. If $s_{\alpha}e_{w}\neq e_{w}$ then $P(e_{w})$ is no $P_{\alpha}$-module, because its $B$-socle has weight $e_{w}$. Thus ${\mathrm{ind}}_{B}^{P_{\alpha}}P(e_{w})={\sf H}_{s_{\alpha}}{\sf H}_{w^{-1}}P(we_{w})$ with $s_{\alpha}\star w^{-1}\neq w^{-1}$ and $s_{\alpha}w^{-1}>w^{-1}$. We have $ws_{\alpha}>w$ in either case, so $w\in W^{P}$. ∎
Lemma 13.9.
Let $\lambda\in X^{P}_{+}$. Let $N\in{\mathrm{rep}}(B)$ be such that all weights $\mu$ of $N$ satisfy $\mu\leq_{a}\lambda$. Then all weights $\mu$ of ${\mathrm{ind}}_{B}^{P}(N)$ satisfy $\mu\leq_{a}\lambda$ and the natural map ${\mathrm{ind}}_{B}^{P}(N)\to N$ induces an isomorphism ${\mathrm{ind}}_{B}^{P}(N)_{\lambda}\to N_{\lambda}$.
Proof.
Recall there is a sequence $\alpha_{1},\cdots,\alpha_{n}$ with $\alpha_{i}\in\Pi^{P}$ and ${\mathrm{ind}}_{B}^{P}={\mathrm{ind}}_{B}^{P_{\alpha_{1}}}\circ\cdots\circ{\mathrm{ind}}_{B}^{P_{\alpha_{n}}}$. Therefore we may assume $P$ is a minimal parabolic, say $P=P_{\alpha}$. Thus $P/B$ is a projective line and we may use [29, II Proposition 5.2]. If $\mu$ is a weight and $(\alpha^{\vee},\mu)\geq 0$, then $s_{\alpha}\mu\leq_{a}\mu$, $\mathrm{R}^{1}{\mathrm{ind}}_{B}^{P}(\mu)=0$, $({\mathrm{ind}}_{B}^{P}(\mu))_{\mu}={\sf k}_{\mu}$ and all weights $\nu$ of ${\mathrm{ind}}_{B}^{P}(\mu)$ lie on the line segment joining $\mu$ with $s_{\alpha}\mu$. In particular, they satisfy $\nu\leq_{a}\mu$. And if $(\alpha^{\vee},\mu)<0$, then ${\mathrm{ind}}_{B}^{P}(\mu)=0$ and all weights $\nu$ of $\mathrm{R}^{1}{\mathrm{ind}}_{B}^{P}(\mu)=0$ lie strictly between the endpoints of the line segment. Thus $\nu<_{a}\mu$. If $N$ has just weight $\lambda$, then it is well known that ${\mathrm{ind}}_{B}^{P}(N)_{\lambda}\to N_{\lambda}$ is an isomorphism [23, Propositions 17, 20]. Now use induction on the number of weights of $N$. ∎
13.1 Generating ${\mathrm{rep}}(P)$

### v3

@@ 12.1 Variations
The ${\mathcal{X}}_{p}$ may depend on the choice of the total order $\prec$ on $W$. Therefore it is not clear that ${\mathcal{X}}_{p}$ is perpendicular to ${\mathcal{X}}_{q}$ when $\ell(p)=\ell(q)$. In fact computer assisted computations in ${\sf K}_{\mathbf{T}}(G/{\mathbf{B}})$ indicate that this fails already for type ${\bf B}_{3}$ with $\ell(p)=\ell(q)=3$ and for type ${\mathbf{C}}_{3}$ with $\ell(p)=\ell(q)=2$.
@@ 13 Generalised flag varieties
Let $\mathbf{P}$ be a parabolic subgroup containing $\mathbf{B}$. (If ${\sf k}={\mathbb{Z}}$, then $\mathbf{P}$ means $\mathbb{P}$.) We seek a full exceptional collection on $\mathbf{G}/\mathbf{P}$. Let $W_{\mathbf{P}}$ be the Weyl group of $\mathbf{P}$, generated inside $W$ by the $s_{\alpha}$ with $\alpha$ simple and $\mathbf{P}_{\alpha}\subset\mathbf{P}$. Let $W^{\mathbf{P}}$ be the set of minimal coset representatives of $W/W_{\mathbf{P}}$, cf.  Lemma 3.7. According to Steinberg [46] we may use as generators of the $R(\mathbf{G})$-module $R(\mathbf{P})$, over ${\mathbb{C}}$, the classes of the irreducible $\mathbf{P}$-modules with highest weight $e_{v}$ where $v$ runs over $W^{\mathbf{P}}$, and $e_{v}$ is still defined as in 4.2. Inspired by that, we restrict our total order $\prec$ from $W$ to $W^{\mathbf{P}}$. Note our convention that the notation $e_{v}$ keeps the meaning it had when $\mathbf{P}=\mathbf{B}$. But only the $e_{v}$ with $v\in W^{\mathbf{P}}$ will be relevant.
@@
The map $W\to W^{\mathbf{P}}$ which sends $w\in W$ to the minimal representative of the coset $wW_{\mathbf{P}}$ is a poset map. If one has chosen a total order on $W^{\mathbf{P}}$ refining the (restriction to $W^{\mathbf{P}}$) of the Bruhat order, then this chosen order can be extended to a total order on $W$ that refines the Bruhat order on $W$. We already had to choose $\prec$ on $W$ in section 3.2, so it makes sense to keep that order $\prec$ and restrict it to $W^{\mathbf{P}}$.
For $v\in W^{\mathbf{P}}$ we will find ${\hat{X}}_{v}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$ such that
@@
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$. Let $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ and $v,w\in W^{\mathbf{P}}$.
@@
    $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes{\hat{X}}_{v},N\otimes{\hat{X}}_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}(M,N)$,
@@
For $v\in W^{\mathbf{P}}$, denote ${\hat{\sf X}}_{v}$ to be the strictly full subcategory of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$ generated by $\{M\otimes{\hat{X}}_{v}\mid M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\}$. Then the category $\mathcal{D}$ has a $\mathbf{G}$-linear semiorthogonal decomposition
$\mathcal{D}=\langle{\hat{\sf X}}_{v}\rangle_{v\in W^{\mathbf{P}}}$ (50)
with respect to the order $\prec$ on $W^{\mathbf{P}}$. Each subcategory ${\hat{\sf X}}_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$.
Definition 13.4.
We put $\hat{\mathcal{X}}_{v}:=\mathcal{L}_{\mathbf{G}/\mathbf{P}}({\hat{X}}_{v})$ for $v\in W^{\mathbf{P}}$.
@@
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{P}))$. Let $v,w\in W^{\mathbf{P}}$.
@@
    The triangulated hull of $\{\hat{\mathcal{X}}_{v}\mid v\in W^{\mathbf{P}}\}$ is $\mathcal{D}$.
In other words, the collection of objects $(\hat{\mathcal{X}}_{v})_{v\in W^{\mathbf{P}}}$ is a full exceptional collection in $\mathcal{D}$.
To prove these Theorems we will need to find replacements–often with hatted notation–of the key ingredients used in the case $\mathbf{P}=\mathbf{B}$. This will take the rest of the section.
Recall that by [32, II Corollary 4.7] we may view ${\mathrm{rep}}(\mathbf{G})$ as a subcategory of ${\mathrm{rep}}(\mathbf{P})$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ as a subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$, so that we may suppress ${\mathrm{res}}^{\mathbf{G}}_{\mathbf{P}}$ in the notation. Similarly we may suppress ${\mathrm{res}}^{\mathbf{P}}_{\mathbf{B}}$. And when we say that a certain $\mathbf{B}$-module $M$ is a $\mathbf{P}$-module, this will mean that $M={\mathrm{res}}^{\mathbf{P}}_{\mathbf{B}}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(M)$.
@@
The strategy for constructing $\hat{\mathcal{X}}_{v}$ is the same as for ${\mathcal{X}}_{v}$, but the construction does not immediately imply the precise relation between ${\mathcal{X}}_{v}$, $\hat{\mathcal{X}}_{v}$ and $\mathbf{G}/\mathbf{B}\to\mathbf{G}/\mathbf{P}$. More specifically, one expects that ${\mathcal{X}}_{v}$ is always the pull back of $\hat{\mathcal{X}}_{v}$ for $v\in W^{\mathbf{P}}$. This is indeed the case:
Theorem 13.7.
Let $v\in W^{\mathbf{P}}$ and let $\pi:\mathbf{G}/\mathbf{B}\to\mathbf{G}/\mathbf{P}$ be the natural map. Then ${\hat{\sf X}}_{v}$ equals ${{\sf X}}_{v}$ in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ and ${\mathcal{X}}_{v}=\pi^{*}\hat{\mathcal{X}}_{v}$.
@@
Notation 13.8.
Let $\Pi^{\mathbf{P}}$ be the set of simple roots $\alpha$ for which $\mathbf{P}_{\alpha}\subset\mathbf{P}$. Let $\Pi^{\notin\mathbf{P}}$ be the set of simple roots outside $\Pi^{\mathbf{P}}$. We say that $\lambda$ is $\mathbf{P}$-dominant if $(\alpha^{\vee},\lambda)\geq 0$ for all $\alpha\in\Pi^{\mathbf{P}}$. The set of $\mathbf{P}$-dominant weights is denoted $X^{\mathbf{P}}_{+}$. Every $W_{\mathbf{P}}$-orbit of weights intersects $X^{\mathbf{P}}_{+}$ in a unique element. Let $\mathbf{G}_{\mathbf{P}}$ be the semisimple subgroup of $\mathbf{P}$ with $\Pi^{\mathbf{P}}$ as simple roots. (So $\mathbf{T}\mathbf{G}_{\mathbf{P}}$ is a Levi subgroup $\mathbf{L}_{\mathbf{P}}$ of $\mathbf{P}$.) Note that $\mathbf{G}_{\mathbf{P}}$ is simply connected.
@@
Let $v\in W^{\mathbf{P}}$. Then $P(-e_{v})$ is a $\mathbf{P}$-module and therefore $P(-e_{v})^{*}$ is a $\mathbf{P}$-module. Conversely, if $e_{w}$ is $\mathbf{P}$-dominant, then $w\in W^{\mathbf{P}}$.
@@
We want to show that $P(-e_{v})={\mathrm{res}}^{\mathbf{P}}_{\mathbf{B}}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(P(-e_{v}))$. Let $\mathbf{P}=\mathbf{U}_{\mathbf{P}}\mathbf{L}_{\mathbf{P}}$ be a Levi decomposition of $\mathbf{P}$ where $\mathbf{U}_{\mathbf{P}}$ is the unipotent radical of $\mathbf{P}$ and $\mathbf{L}_{\mathbf{P}}$ is the Levi component. Let $w_{0}^{\mathbf{P}}\in W_{\mathbf{P}}$ be the longest element and consider a reduced decomposition $w_{0}^{\mathbf{P}}=s_{1}\cdots s_{l}$ of $w_{0}^{\mathbf{P}}$. The functors ${\sf H}_{w_{0}^{\mathbf{P}}}$ and ${\mathrm{res}}^{\mathbf{P}}_{\mathbf{B}}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}$ are identical, because $\overline{\mathbf{B}w_{0}^{\mathbf{P}}\mathbf{B}}/\mathbf{B}=\mathbf{P}/\mathbf{B}$. So we now want to prove that ${\sf H}_{w_{0}^{\mathbf{P}}}(P(-e_{v}))=P(-e_{v})$. By Lemma 3.1 we have ${\sf H}_{w_{0}^{\mathbf{P}}}={\sf H}_{s_{1}}\circ\cdots\circ{\sf H}_{s_{l}}$. So it suffices to show that ${\sf H}_{s}(P(-e_{v}))=P(-e_{v})$ for $s\in W_{\mathbf{P}}$ simple. Now $v$ is a minimal coset representative in $W/W_{\mathbf{P}}$, so $vs>v$ for $s\in W_{\mathbf{P}}$. Then $sv^{-1}>v^{-1}$, so $sv^{-1}w_{0}<v^{-1}w_{0}$, and $s\star v^{-1}w_{0}=v^{-1}w_{0}$. Thus ${\sf H}_{s}(P(-e_{v}))={{\sf H}}_{s}{{\sf H}}_{v^{-1}w_{0}}(-w_{0}ve_{v})={{\sf H}}_{v^{-1}w_{0}}(-w_{0}ve_{v})=P(-e_{v})$ (Cf. Lemma 3.1).
Conversely, let $e_{w}$ be $\mathbf{P}$-dominant and $\alpha\in\Pi^{\mathbf{P}}$. Now $we_{w}$ is dominant and $w^{-1}$ is minimal amongst the $z$ with $zwe_{w}=e_{w}$. Thus if $s_{\alpha}e_{w}=e_{w}$ then $s_{\alpha}w^{-1}>w^{-1}$. If $s_{\alpha}e_{w}\neq e_{w}$ then $P(e_{w})$ is no $\mathbf{P}_{\alpha}$-module, because its $\mathbf{B}$-socle has weight $e_{w}$. Thus ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}_{\alpha}}P(e_{w})={\sf H}_{s_{\alpha}}{\sf H}_{w^{-1}}P(we_{w})$ with $s_{\alpha}\star w^{-1}\neq w^{-1}$ and $s_{\alpha}w^{-1}>w^{-1}$. We have $ws_{\alpha}>w$ in either case, so $w\in W^{\mathbf{P}}$.
@@
Let $\lambda\in X^{\mathbf{P}}_{+}$. Let $N\in{\mathrm{rep}}(\mathbf{B})$ be such that all weights $\mu$ of $N$ satisfy $\mu\leq_{a}\lambda$. Then all weights $\mu$ of ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(N)$ satisfy $\mu\leq_{a}\lambda$ and the natural map ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(N)\to N$ induces an isomorphism ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(N)_{\lambda}\to N_{\lambda}$.
@@
Recall there is a a reduced decomposition of $w_{0}^{\mathbf{P}}=s_{1}\cdots s_{l}$ of $w_{0}^{\mathbf{P}}$ and that ${\mathrm{res}}^{\mathbf{P}}_{\mathbf{B}}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}={\sf H}_{s_{1}}\circ\cdots\circ{\sf H}_{s_{l}}$. Therefore we may assume $\mathbf{P}$ is a minimal parabolic, say $\mathbf{P}=\mathbf{P}_{\alpha}$. Thus $\mathbf{P}/\mathbf{B}$ is a projective line and we may use [32, II Proposition 5.2]. If $\mu$ is a weight and $(\alpha^{\vee},\mu)\geq 0$, then $s_{\alpha}\mu\leq_{a}\mu$, $\mathrm{R}^{1}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(\mu)=0$, $({\mathrm{ind}}_{\mathbf{B}}^{P}(\mu))_{\mu}={\sf k}_{\mu}$ and all weights $\nu$ of ${\mathrm{ind}}_{\mathbf{B}}^{P}(\mu)$ lie on the line segment joining $\mu$ with $s_{\alpha}\mu$. In particular, they satisfy $\nu\leq_{a}\mu$. And if $(\alpha^{\vee},\mu)<0$, then ${\mathrm{ind}}_{\mathbf{B}}^{P}(\mu)=0$ and all weights $\nu$ of $\mathrm{R}^{1}{\mathrm{ind}}_{\mathbf{B}}^{P}(\mu)$ lie strictly between the endpoints of the line segment. Thus $\nu<_{a}\mu$. If $N$ has just weight $\lambda$, then it is well known that ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(N)_{\lambda}\to N_{\lambda}$ is an isomorphism [26, Propositions 17, 20]. Now use induction on the number of weights of $N$. ∎
13.1 Generating ${\mathrm{rep}}(\mathbf{P})$